OpenStudy (shamim):

if A and B are the two vectors representing two diagonals of a parallelogram. what is the area of that parallelogram

OpenStudy (anonymous):

One diagonal would be the vector sum of adjacent sides .... A=a+b The other diagonal would be the vector difference between adjacent sides B=a-b A+B = 2a .... a=(A+B)/2 A-B =2b .... b=(A-B)/2 a cross b = area